OpenStudy (anonymous):

Solve: -8√(5x-5)=-3 A: -5 B: 5 C: 6 D: -6

OpenStudy (anonymous):

OH! I figured it out, I just forgot to put the + sign after the -8, that's why I was confused.

OpenStudy (australopithecus):

so you know how to deal with radicals then?

OpenStudy (anonymous):

Yup, just made a silly math error. Thanks for your help, though!

OpenStudy (imtiaz7):

first take the -8 under the 3 and then squrae on both side in this you can solve easly.

OpenStudy (australopithecus):

First realize that, \[\sqrt{x} = x^{\frac{1}{2}}\] so we can write your problem as follows, \[-8(5x-5)^{\frac{1}{2}} = -3\] you need to eliminate the radical to solve for x, to do this you need to raise both sides of the equation by a certain power that will give you a 1 or a whole number. For example \[x^{\frac{1}{3}} = 12\] can be solved by raising both sides to the power 3 \[(x^{\frac{1}{3}})^3 = 12^3\] \[x^{\frac{1}{3}*3} = 12^3\] \[x^{\frac{1*3}{3}} = 12^3\] \[x^{\frac{3}{3}} = x^1 = x = 12^3\]